# Leveraging Multi-lingual Positive Instances in Contrastive Learning to Improve Sentence Embedding

### v1

@@ 3.5 Main Experimental Results
More specifically, we aim to address two questions based on the overall results of various downstream tasks shown in Table 1. Q1: Does the utilization of multiple positives yield more substantial benefits to the model compared to conventional CL with a single positive? Q2: Does the effectiveness of leveraging multiple positives still hold for stronger sentence embedding models?
3.5.1 Leveraging Multiple Positives Helps Learning Multi-lingual Sentence Embedding with CL.
@@
To compare their performance, we first train general language models with a warm-up under our proposed methods and conventional CL. Specifically, we choose BERT-base-multilingual-uncased  [^3] https://huggingface.co/bert-base-multilingual-uncased and XLM-RoBERTa-large  [^4] https://huggingface.co/xlm-roberta-large.
@@
From the upper part of Table 1, it is evident that continuous training with both single and multiple positives on pretrained language models enhances the performance of the models, making them to competitive level. However, from the average score of the downstream tasks, we can observe that models trained using our proposed framework with multiple positives demonstrate stronger performance than conventional CL-based methods with single positives. Specifically, we discover an average improvement of 1.0 for mBERT and 1.5 for XLM-R across three different downstream tasks compared to models trained on single positives. The most significant improvement is seen in STS for mBERT, and in MTOP Domain Classification for XLM-R.
In order to fairly compare MPCL with state-of-the-art models, we train a variant with the same language coverage using hard negatives. We include all 15 languages in XNLI and utilize sentences with contradictory labels as hard negatives. This comparison is referred to as XLM-R w/ hard negative shown in Table 1. Our XLM-R trained with multiple positives surpasses mSimCSE, which is trained with single positives in all tasks we report. Note that during the training of this variant, an anchor sentence can access five positives from different languages and the hard negative instance is also randomly selected from all languages. This variant indicates that our proposed MPCL can be incorporated with other existing orthogonal methods, such as hard negatives, and contributes to better performance.
@@
These observations suggest that for under-exploited multi-lingual language models, leveraging multiple positives can offer a richer and more useful source of information for training, thus yielding more substantial benefits to the model. Note that with the exception of the BUCC task, all the other evaluation tasks include languages that are excluded from our dataset.
3.5.2 Leveraging Multiple Positives Can Further Improve Sentence Embedding Models.
@@
Next, we continually train strong backbones, including LaBSE and mSimCSE to validate the effectiveness of multiple positives for pretrained sentence embedding models. Their results of BUCC and Tatoeba are adopted from Wang et al. (2022b). For mSimCSEall, we use its released checkpoint  [^5] https://github.com/yaushian/mSimCSE to evaluate its performance on STS17, STS22, and MTOP Domain Classification. LaBSE’s results are adopted from the MTEB leaderboard.
@@
Upon examining the lower part of Table 1, we can observe that the highest scores for each task are obtained by models continually trained on our framework. At the same time, it is evident that the improvements in the BUCC, STS, and MTOP Domain Classification tasks are more prominent compared to the Tatoeba dataset. A possible reason for this observation is that we only include five languages, excluding English in our dataset, in contrast to the comprehensive evaluation of 36 languages in the Tatoeba dataset. These findings imply that the utilization of multiple positives can even significantly enhance the performance of well-trained sentence embedding models, indicating the robustness of our proposed method. Refer to Appendix A.2 for detailed results.
@@
Figure 3: Detailed score changes of LaBSE and mSimCSEall on different languages after training on single positive and multiple positives. The horizontal axis 0 stands for the original scores of models. The solid lines splits the results into different part: STS17, STS22 and MTOP Domain Classification from left to right. The dash squares highlight unseen languages and their average performance for each particular task.
3.6 Transfer Performance on Unseen Languages.
@@
The horizontal axis 0 stands for the original scores of LaBSE and mSimCSEall. For example, in Figure 3, the first pair (en-de) indicates that parallel training with a single positive leads to a slight decrease in performance for both LaBSE and mSimCSE, while training with multiple positives leads to an improvement. Note that our training dataset only contains en, de, fr, es, ru and zh so all the other language pairs are considered unseen. Language pairs such as en-tr, where only one of the two languages is observed in the training set are also considered as exclusive.
From the right part of Figure 3, where we report the average change on three tasks, it is evident that models trained with multiple positives generally exhibit better transfer ability to unseen languages. For example, when looking at the unseen language tr in STS17 and STS22, we observe a drop in accuracy with single positive training, while multiple positive training still shows a significant improvement in LaBSE. This trend is similarly observed in mSimCSE, indicating that the transfer ability of multiple positives surpasses that of single positive training.
3.7 The Choice of Languages Included in the Training Dataset.
@@
3.7.1 Training with Only Unseen Language.
We first construct another training dataset using multiple positives extracted from XNLI which comprises only four languages (bg, el, vi and sw) that do not overlap with any language that will be tested in STS17 and STS22 so that all results exhibit models’ fully transfer abilities. Besides, we add two more languages (hi and th) to see whether cross-lingual signals from more languages that do not overlap with test languages can help improve the transfer ability. Figure 4 shows the results.
Figure 4: Results of the STS17 and STS22 tasks trained on fully none-overlapping languages. We report the average scores of all language pairs included in STS17 and STS22.
As reported in Wang et al. (2022b), we also observe that CL-based methods exhibit remarkable transfer abilities to totally unseen languages. From Figure 4, we can observe a slight drop when incorporating two additional languages into the dataset. This finding aligns with the observation presented by Conneau et al. (2020), where they highlight a trade-off between the number of languages and transfer performance. However, the strong transferability of multiple positives remains evident as training on completely non-overlapping languages can still yield competitive results compared to our original training dataset. Besides, we also observe an obvious trend in Figure 4 where all average results from multiple positives consistently surpass those from single positives. We believe that by bringing anchor sentences and all the remaining positives closer together, models can effectively capture the nuances of different languages and achieve better language representation.
3.7.2 Training with Only Seen Language.
We also explore the effect of training with only languages that will be tested in the STS17 and STS22 tasks. We remove the ru and zh from our dataset, which will not be evaluated in STS17 to see whether ru and zh provide more useful cross-lingual signals. Besides, we add two more additional languages, ar, and tr so that all the languages in this dataset (en, de, fr, es, ar, tr) will be tested in the STS17 and STS22 tasks. The results are shown in Table 2.
@@
   +Single 67.9 68.9 71.1 57.2 59.1 64.8 65.6 67.1 60.0
   +Multiple 71.2 71.9 72.9 57.1 60.2 64.1 64.8 68.6 61.3
   +Single, -ru, zh 67.7 71.4 71.4 55.2 57.3 63.6 65.1 67.0 59.2
   +Multiple, -ru, zh 70.5 71.1 72.0 56.2 57.5 62.4 65.7 67.8 59.9
   +Single, +ar, tr 71.9 72.3 71.5 58.6 58.7 66.0 64.4 66.8 60.4
   +Multiple, +ar, tr 72.6 74.3 72.0 59.2 58.9 65.3 66.4 66.9 60.9
Table 2: Results of different composition of dataset on STS17 and STS22 task. ru,zh-ru,zh means we only have four languages in the dataset while +ar,tr+ar,tr means that we replace ru and zh with ar and tr.
We focus specifically on the language pairs that we removed or added such as en-ar and en-tr to examine the impact of different dataset compositions. Upon removing ru and zh from the original dataset, we observe a slight decrease in the overall accuracy of STS17. This suggests that ru and zh in fact contribute valuable cross-lingual signals to the model in STS17. However, in STS22, although the performance of ru drops due to its removal, the accuracy of zh acquires an interesting improvement. One possible reason for this may be that although we include zh in our dataset, we do not have the exact same zh-zh pairs that are tested in STS22. Thus, the model is able to learn information from other cross-lingual signals.
@@ 4.1 Contrastive Learning
In addition, some researchers have already applied contrastive learning with multiple positives and showed the effectiveness of multiple positives Hoffmann et al. (2022); Khosla et al. (2020). Dwibedi et al. (2021) and Frosst et al. (2019) have demonstrated that incorporating more positive instances in contrastive learning significantly improves the models’ performance.
@@ 4.2 Mono-lingual Sentence Embeddings
Gao et al. (2021) propose SimCSE exploits the ability of CL by using dropout as augmentation in unsupervised settings. For supervised sentence embeddings, SimCSE makes use of the NLI dataset Bowman et al. (2015) to establish positive and hard negative examples. The success of SimCSE attracts researchers’ attention to CL when dealing with sentence embeddings Ni et al. (2022); Wang et al. (2022a); Su et al. (2022); Xie et al. (2022) since CL-based models provide more competitive results on downstream tasks than classical models such as S-BERT (Sentence-BERT) Reimers and Gurevych (2019). Su et al. (2022) combine CL with prompt-like instructions and propose INSTRUCTOR, which encodes input sentences and instructions for a specific task together to help the model understand what exact task it is going to do.
@@ 4.3 Multi-lingual Sentence Embeddings
Training a universal sentence embedding model for all languages is a fundamental and important task. As the ability of efficient similarity calculation across languages, multi-lingual sentence embeddings have been applied to low-resource corpus filtering Chaudhary et al. (2019), parallel corpus mining Kvapilıková et al. (2021), and synthesized dataset filtering Wu et al. (2023). Artetxe and Schwenk (2019) utilize the BiLSTM structure with a shared vocabulary for all languages. Recent studies prefer to adopt CL in multilingual settings with source-target translation pairs where the target sentence would be considered as the particular positive for the source sentence. For example, LaBSE Feng et al. (2020) is a BERT-based multilingual sentence representation model trained on massive amounts of monolingual data and translation pairs covering over 100 languages. Wang et al. (2022b) propose a multilingual version of SimCSE, called mSimCSE, showing that the contrastive learning objective of SimCSE can also be applied to multi-lingual settings by using multi-lingual data. Although CL-based models have become common, leveraging multiple positives for learning multi-lingual sentence embeddings is still underexplored.
@@ 5 Conclusion
Although we explore the effect of using multi-lingual translations as multiple positives, our experiments are restricted by the limited amount of multi-lingual translations. Even the XNLI data we used are machine-translated, not human-translated high-quality ones. Besides, our choice of training languages is also restricted by the XNLI dataset. The composition of the datasets and how one positive language can influence another remains to be explored.
@@ A.1 Examples we use in Figure 1
For mono-lingual examples, we randomly sample 100 sentences with entailments from the English XNLI dataset  [^7] https://huggingface.co/datasets/xnli. The premise is used as anchor and the corresponding entailment sentence serves as one of the positives. Other positives are generated by ChatGPT and the prompt we used is “Give me several sentences that share similar meaning with the following one: ”. For multilingual examples, we use the same 100 sentences and their corresponding translations from the XNLI dataset and calculate their average similarity scores. We report the average similarity score of these sentences in Figure 1 (c) and (d).
Model fr ru zh de avg.
mSimCSEall - - - - 95.2
   +Single 94.9 96.4 96.8 96.3 96.1
   +Multiple 95.2 96.9 96.4 96.9 96.4
LaBSE - - - - 93.5
   +Single 96.4 97.6 97.0 98.2 97.3
   +Multiple 97.1 97.8 97.8 98.1 97.7
Table 3: Full results of BUCC task.
@@ A.2 Detailed Results of Main Experiments
Model de fr es ru zh avgin tr hi 14 avg. 36 avg.
mSimCSEall (Reproduced) 98.8 94.7 97.2 94.2 96.5 96.3 95.8 96.2 93.2 91.2
   + Single 98.8 95.6 98.1 93.8 96.4 96.5 96.5 96.8 93.3 91.2
   + Multiple 99.0 95.2 98.2 94.7 96.0 96.6 96.8 97.0 93.5 91.3
LaBSE (Reproduced) 99.0 96.0 96.3 95.3 96.1 96.5 98.3 97.3 95.8 95.3
   + Single 99.2 96.0 98.4 95.0 96.1 96.9 98.4 97.8 95.7 95.1
   + Multiple 99.4 96.4 98.3 95.2 96.4 97.1 98.5 97.9 96.1 95.5
Table 6: Detailed results of Tatoeba dataset. Since in the paper of LaBSE Feng et al. (2020) and mSimCSE Wang et al. (2022b), authors did not report the score for each language, we reproduce their scores on Tatoeba dataset through XTREME benckmark to have a better comparison. avgin stands for the languages included in our training dataset. tr and hi are languages excluded from our training dataset.
@@
   + Multiple 78.0 76.1 77.0 62.6 59.9 61.3
Table 7: Results of STS tasks. We evaluate mSimCSEall through MTEB by ourselves since authors did not do this experiment. Taskin stands for language pairs, both multi-lingual and cross-lingual, that are inside our training set while Taskex stands for exclusive language pairs. Language pairs such as en-tr, where only one of the two languages is observed in training set are also considered as exclusive. More specifically, in STS17, there are five included pairs and six excluded pairs while in STS22, there are ten and seven, respectively (with no fr-pl).
@@
Table 8: Full results of MTOP domain classification. Notice that hi and th are languages excluded from our training dataset.

### v2

@@ 3.5 Main Experimental Results
More specifically, we aim to address two questions based on the overall results of various downstream tasks shown in Table 1. Q1: Does the utilization of multiple positives yield more substantial benefits to the model compared to conventional CL with a single positive? Q2: Does the effectiveness of leveraging multiple positives still hold for stronger sentence embedding models? Refer to Appendix A.2 for detailed results of different models.
3.5.1 Multiple Positives Yield Better Performance than Single Positive
@@
We first train general LMs with a warm-up under our proposed methods and compare their performance with LMs trained on conventional CL.
@@
From the upper part of Table 1, it is evident that continuous training with both single and multiple positives on pretrained language models enhances the performance of the models. However, from the average score of the downstream tasks, we can observe that models trained using our proposed framework with multiple positives demonstrate stronger performance than conventional CL-based methods with single positives. Specifically, we discover an average improvement of 1.1 for mBERT and 1.3 for XLM-R across different downstream tasks compared to models trained on single positive. The most significant improvement is seen in STS22 for mBERT, and in STS17 for XLM-R.
[…]
@@
These observations suggest that for general multi-lingual language models, leveraging multiple positives can offer a richer and more useful source of information for training, thus yielding more substantial benefits to the model. Note that with the exception of the BUCC task, all the other evaluation tasks include languages that are excluded from our dataset.
3.5.2 Multiple Positives Can Further Improve Sentence Embedding Models
@@
Next, we continually train sentence embedding models, including LaBSE and mSimCSE to validate the effectiveness of multiple positives for pretrained sentence embedding models. For mSimCSEall, their results of BUCC and Tatoeba are adopted from Wang et al. (2022b) while we use its released checkpoint  [^5] https://github.com/yaushian/mSimCSE to evaluate its performance on STS17, STS22, and MTOP Domain Classification. LaBSE’s results are adopted from the MTEB leaderboard.
@@
Upon examining the lower part of Table 1, we can observe that the highest scores for each task are obtained by models continually trained on our framework. At the same time, it is evident that the improvements in the BUCC, STS, and MTOP Domain Classification tasks are more prominent compared to the Tatoeba dataset. A possible reason for this observation is that we only include five languages, excluding English in our dataset, in contrast to the comprehensive evaluation of 36 languages in the Tatoeba dataset. These findings imply that the utilization of multiple positives can even significantly enhance the performance of well-trained sentence embedding models, indicating the robustness of our proposed method.
(a) Performance variation on LaBSE
(b) Performance variation on mSimCSEall
@@
Figure 3: Detailed performance changes of LaBSE and mSimCSEall on different languages after training on multiple positive or single positive. On the horizontal axis, 0 represents the original scores without additional training. The vertical axis shows changes in performance after further training. Bars above the horizontal axis indicate improvements, while those below indicate decreases. The bold lines split the results into three different parts: STS17, STS22 and MTOP Domain Classification from left to right. Orange color highlights unseen languages.
3.6 Transferring to Unseen Languages
@@
The value of 0 on horizontal axis stands for the scores of the origianl LaBSE and mSimCSEall. For example, in Figure 3, the first pair English-German (en-de) indicates that parallel training with a single positive leads to a slight decrease in performance for both LaBSE and mSimCSE, while training with multiple positives leads to an improvement. Note that our training dataset only contains en, de, fr, es, ru and zh so all the other language pairs are considered unseen. Language pairs such as English-Turkish (en-tr), where only one of the two languages is observed in the training set are also considered as exclusive.
In Figure 3, we can observe that models trained with multiple positives generally exhibit better transfer ability to unseen languages. For example, when looking at the unseen language pair pl-en in STS22, we observe a drop in accuracy with single positive training, while multiple positive training still shows a significant improvement in LaBSE. This trend can also be observed in mSimCSE’s variance, indicating that the transfer ability of multiple positives surpasses that of single positive training. Average results of unseen languages in STS tasks can be found in Appendix A.2, Table 7.
3.7 The Choice of Languages in Training Set.
@@
3.7.1 Training with Only Unseen Language
We first construct another training dataset using multiple positives extracted from XNLI which comprises only four languages: Bulgarian (bg), Greek (el), Vietnamese (vi) and Swahili (sw) that do not overlap with any language that will be tested in STS17 and STS22 so that all results exhibit models’ fully transfer abilities. Besides, we add two more languages: Hindi (hi) and Thai (th) to see whether cross-lingual signals from more non-overlap languages can help improve the transfer ability. Figure 4 shows the results.
Figure 4: Results of the STS17 and STS22 tasks trained on fully none-overlapping languages. We report the average scores of all language pairs included in STS17 and STS22. Error bars refer to standard deviation.
As reported in Wang et al. (2022b), we also observe that CL-based methods exhibit remarkable transfer abilities on totally unseen languages. From Figure 4, we can observe a slight drop when incorporating two additional languages into the dataset. This finding aligns with the observation presented by Conneau et al. (2020), where they highlight a trade-off between the number of languages and transfer performance. However, the strong transfer ability of multiple positives remains evident as training on completely non-overlapping languages can still yield competitive results compared to our original training dataset. Besides, we also observe an obvious trend in Figure 4 where all average results from multiple positives consistently surpass those from single positives. We believe that by bringing anchor sentences and all the remaining positives closer together, models can effectively capture the nuances of different languages and achieve better language representation.
3.7.2 Training with Only Seen Language
We also explore the effect of training with only languages that will be tested in the STS17 and STS22 tasks. We remove the ru and zh from our dataset, which will not be evaluated in STS17 to see whether ru and zh provide more useful cross-lingual signals. Besides, we add two more additional languages, Arabic (ar), and Turkish (tr) so that all the languages in this dataset (en, de, fr, es, ar, tr) will be tested in the STS17 and STS22 tasks. The results are shown in Table 2.
@@ 3.7.2 Training with Only Seen Language
   +Single 67.9±0.49 68.4±0.71 71.1±0.63 57.2±0.64 59.1±0.39 64.4±0.43 65.6±0.25 67.4±0.65 59.8±0.44
   +Multiple 71.0±0.24 72.5±0.40 73.2±0.20 57.1±0.48 60.3±0.20 64.3±0.23 64.8±0.27 68.6±0.22 61.4±0.41
   +Single, -ru, zh 67.6±0.18 70.3±0.88 70.6±0.64 55.6±0.46 57.5±0.28 63.1±0.60 65.3±0.17 67.3±0.42 59.6±0.46
   +Multiple, -ru, zh 70.3±0.49 70.8±0.23 71.8±0.48 56.2±0.07 58.4±0.67 62.7±0.60 65.7±0.28 68.1±0.32 60.2±0.31
   +Single, +ar, tr 71.2±0.61 72.1±0.58 71.3±0.44 58.7±0.33 58.3±0.32 65.7±0.26 64.8±0.51 67.0±0.37 60.3±0.36
   +Multiple, +ar, tr 72.5±0.18 73.6±0.49 72.1±0.27 59.0±0.27 59.3±0.61 65.3±0.41 66.3±0.34 67.3±0.35 60.9±0.34
Table 2: Results of different composition of dataset on STS17 and STS22 task. -ru, zh means we only have four languages in the dataset while +ar, tr means that we replace ru and zh with ar and tr.
We focus specifically on the language pairs that we removed or added such as English-Arabic (en-ar) and English-Turkish (en-tr) to examine the impact of different dataset compositions. Upon removing ru and zh from the original dataset, we observe a slight decrease in the overall accuracy of STS17. This suggests that ru and zh in fact contribute valuable cross-lingual signals to the model in STS17. However, in STS22, although the performance of ru drops due to its removal, the accuracy of zh acquires an interesting improvement. One possible reason for this may be that although we include zh in our dataset, we do not have the exact same zh-zh pairs that are tested in STS22. Thus, the model is able to learn information from other cross-lingual signals.
@@
3.8 Case study
In this section, we randomly choose two examples from XNLI test set to demonstrate effect of multiple positives on cross-lingual similarity. With the six languages included in our experiment, we can obtain fifteen cross-lingual pairs. The similarities are calculated for all fifteen language pairs. The results are shown in Figure 5. The sub-figure caption is the example sentence in English. Non-English-centric pairs are highlighted with red color. As we calculate similarity scores based on translations, the gold label similarity score for all cross-lingual pairs should be 1.0. As shown in the figure, it can be observed that for most English-centric pairs, single positive and multiple positive models achieve a comparable level of similarity. However, for non-English-centric pairs, multiple positive model exhibits an obvious higher similarity. This indicates that our approach of utilizing translations as multiple positives improves the cross-lingual representation learning, especially for non-English-centric language pairs.
@@ 4.1 Contrastive Learning
Contrastive learning with multiple positives has been studied in previous researches in computer vision Khosla et al. (2020), with fine-grained strategies such as soft-nearest neighbor Frosst et al. (2019) and ranking Dwibedi et al. (2021); Hoffmann et al. (2022). In this paper, as the focus is verifying the impact of the usage of multiple translated positives for sentence embedding, we simply assign equal importance to all positives.
(a) “We had a great talk.”
(b) “I don’t know whether he stayed in Augusta after that.”
Figure 5: Cross-lingual similarity scores calculated by XLM-R + Multiple or XLM-R + Single for two randomly chosen examples. The example sentences are shown in the sub-figure caption. Non-English-centric language pairs are highlighted with red color.
@@ 4.2 Mono-lingual Sentence Embeddings
SimCSE Gao et al. (2021) exploits the ability of CL by using dropout as augmentation in unsupervised settings. For supervised sentence embeddings, SimCSE makes use of the NLI dataset Bowman et al. (2015) to establish positive and hard negative examples. The success of SimCSE attracts researchers’ attention to CL when dealing with sentence embeddings Ni et al. (2022); Wang et al. (2022a); Su et al. (2022); Xie et al. (2022) since CL-based models provide more competitive results on downstream tasks than classical models such as S-BERT (Sentence-BERT) Reimers and Gurevych (2019). Su et al. (2022) further combine CL with prompt-like instructions while Liu et al. (2023) leverages ranking information into CL.
@@ 4.3 Multi-lingual Sentence Embeddings
Training a universal sentence embedding model for all languages is a fundamental and important task. As the ability of efficient similarity calculation across languages, multi-lingual sentence embeddings have been applied to low-resource corpus filtering Chaudhary et al. (2019), parallel corpus mining Kvapilıková et al. (2021), and synthesized dataset filtering Wu et al. (2023). Artetxe and Schwenk (2019) utilize the BiLSTM structure with a shared vocabulary for all languages. Recent studies prefer to adopt CL in multilingual settings with source-target translation pairs where the target sentence would be considered as the particular positive for the source sentence. For example, LaBSE Feng et al. (2020) is a BERT-based multilingual sentence representation model trained on massive amounts of monolingual data and translation pairs covering over 100 languages. Mao and Nakagawa (2023) distill LaBSE for lightweight mutil-lingual sentence embedding models. Wang et al. (2022b) show that CL resembling SimCSE can also be applied to multi-lingual settings by using multi-lingual data. Besides, sentence-level CL is often combined with token-level information, for example, token-level reconstruction Mao et al. (2022) and token-level alignment Li et al. (2023) to improve cross-lingual sentence embeddings. Although CL-based models have become common, leveraging multiple positives for learning multi-lingual sentence embeddings is still unexplored.
@@ 5 Conclusion
Although we explore the effect of using multi-lingual translations as multiple positives, the experiments are still limited by the number of languages. Study on more low-resource languages could be taken into consideration. Besides, the XNLI data we used are machine-translated, with possible noises within. The composition of the training languages and how the training languages can affect the testing on other languages also remains to be explored. Additionally, as we study the impact of multiple multi-lingual positives for sentence embedding in this paper, the positives are assigned with equal importance. But there are various fine-grained strategies to weight positives such as ranking deserving exploited.
@@ A.1 Details of Figure 1
For mono-lingual examples, we randomly sample 100 sentences with entailments from the English XNLI dataset  [^7] https://huggingface.co/datasets/xnli. The premise is used as anchor and the corresponding entailment sentence serves as one of the positives. Other positives are generated by ChatGPT  [^8] https://chat.openai.com/ and the prompt we used is “Give me several sentences that share similar meaning with the following one: ”. For multilingual examples, we use the same 100 sentences and their corresponding translations from the XNLI dataset and calculate their similarity scores. We report the average similarity score of these sentences in Figure 1 (c) and (d).
@@ A.2 Detailed Results of Main Experiments
Model fr ru zh de avg.
mBERT - - - - 56.7
   + Single 85.2 83.1 80.2 88.3 84.1
   + Multiple w/o hard negative 86.7 84.2 81.0 89.3 85.3
XLM-R - - - - 66.0
   + Single 94.2 95.1 93.1 95.2 94.5
   + Multiple w/o hard negative 94.9 96.0 95.2 96.4 95.7
   + Multiple w/ hard negative 94.5 95.0 96.6 95.5 95.4
mSimCSEall - - - - 95.2
   +Single 94.9 96.4 96.8 96.3 96.1
   +Multiple 95.1 96.5 96.2 96.4 96.0
LaBSE - - - - 93.5
   +Single 96.4 97.6 97.0 98.2 97.3
   +Multiple 96.9 97.8 97.6 98.1 97.6
Table 5: Full results of BUCC task.
Model de fr es ru zh avgin 14 avg. 36 avg.
mBERT - - - - - - - -
   + Single 95.8 89.8 88.4 86.3 87.2 89.5 70.5 64.4
   + Multiple w/o hard negative 96.2 89.8 90.1 87.9 89.1 90.6 71.6 65.1
XLM-R - - - - - - 57.6 53.4
   + Single 98.6 95.0 97.6 93.6 95.6 96.1 91.3 89.6
   + Multiple w/o hard negative 99.0 95.2 97.8 94.1 96.5 96.5 92.0 90.4
   + Multiple w/ hard negative 98.9 95.6 97.8 94.4 96.2 96.6 93.5 91.8
mSimCSEall (Reproduced) 98.8 94.7 97.2 94.2 96.5 96.3 93.2 91.2
   + Single 98.8 95.6 98.1 93.8 96.4 96.5 93.3 91.2
   + Multiple 99.0 95.2 98.2 94.7 96.0 96.6 93.5 91.4
LaBSE (Reproduced) 99.0 96.0 96.3 95.3 96.1 96.5 95.3 95.0
   + Single 99.2 96.0 98.4 95.0 96.1 96.9 95.7 95.1
   + Multiple 99.4 96.4 98.3 95.2 96.4 97.1 96.0 95.4
Table 6: Detailed results of Tatoeba dataset. Since in the paper of LaBSE Feng et al. (2020) and mSimCSE Wang et al. (2022b), authors did not report the score for each language, we reproduce their scores on Tatoeba dataset through XTREME benchmark to have a better comparison. avgin stands for the languages included in our training dataset.
@@
mBERT - - - - - -
   + Single 63.7 51.5 57.0 54.3 52.2 53.4
   + Multiple 63.8 52.8 57.8 56.8 54.5 55.8
XLM-R - - - - - -
   + Single 73.8 68.9 71.1 61.2 57.8 59.8
   + Multiple w/o hard negative 74.8 71.9 73.2 63.2 58.9 61.4
   + Multiple w/ hard negative 81.1 77.1 78.9 65.1 62.5 64.0
@@
   + Multiple 78.0 76.1 76.9 62.6 59.9 61.5
Table 7: Detailed results of STS tasks. We evaluate mSimCSEall through MTEB by ourselves. Taskin stands for language pairs, that are inside our training set while Taskex stands for exclusive language pairs. More specifically, in STS17, there are five included pairs and six excluded pairs while in STS22, there are ten and seven, respectively (excluding fr-pl).
@@
mBERT - - - - - - -
   + Single 74.6 78.0 75.1 69.3 59.8 16.9 62.3
   + Multiple 74.1 78.6 76.0 69.9 60.5 16.8 62.7
XLM-R - - - - - - -
   + Single 84.5 85.4 85.0 81.7 79.2 82.1 83.0
   + Multiple w/o hard negative 86.3 86.2 86.8 82.3 83.1 82.0 84.5
   + Multiple w/ hard negative 88.0 88.5 88.8 86.2 85.9 83.8 86.8
@@
Table 8: Full results of MTOP domain classification. We report the accuracy metric on test set. Notice that hi and th are languages excluded from our training dataset.